\begin{thebibliography}{99}

\bibitem{AD90}
Sergei Artemov and Giorgie Dzhaparidze.
\newblock Finite {K}ripke models and predicate logics of provability.
\newblock {\em The Journal of Symbolic Logic}, 55(3):1090--1098, 1990.

\bibitem{Behmann22}
H.~Behmann.
\newblock Beitr\"{a}ige z\"{u}r {A}lgebra der {L}ogik, inbesondere zum
  {E}ntscheidungsproblem.
\newblock {\em Mathematische Annalen}, 86:163--229, 1922.

\bibitem{Berger66}
Robert Berger.
\newblock {\em The {U}ndecidability of the {D}omino {P}roblem}, volume~66 of
  {\em Memoirs of {A}{M}{S}}.
\newblock {A}{M}{S}, 1966.

\bibitem{BGG97}
Egon B\"{o}rger, Erich Gr\"adel, and Yuri Gurevich.
\newblock {\em The Classical Decision Problem}.
\newblock Springer, 1997.

\bibitem{ChRyb03}
Alexander Chagrov and Mikhail Rybakov.
\newblock How many variables does one need to prove {P}{S}{P}{A}{C}{E}-hardness
  of modal logics?
\newblock In Philippe Balbiani, Nobu-Yuki Suzuki, Frank Wolter, and Michael
  Zakharyaschev, editors, {\em Advances in Modal Logic 4}, pages 71--82. King's
  College Publications, 2003.

\bibitem{Corsi93}
Giovanna Corsi.
\newblock Quantified modal logics of positive rational numbers and some related
  systems.
\newblock {\em Notre Dame Journal of Formal Logic}, 34(2):263--283, 1993.

\bibitem{FM98}
Melvin Fitting and Richard~L. Mendelsohn.
\newblock {\em First-{O}rder Modal Logic}, volume 277 of {\em Synthese
  Library}.
\newblock Kluwer Academic Publishers, 1998.

\bibitem{Gabbay81}
Dov Gabbay.
\newblock {\em Semantical Investigations in Heyting's Intuitionistic Logic}.
\newblock D. Reidel, 1981.

\bibitem{GKWZ}
Dov Gabbay, Agi Kurucz, Frank Wolter, and Michael Zakharyaschev.
\newblock {\em Many-{D}imensional Modal Logics: Theory and Applications},
  volume 148 of {\em Studies in Logic and the Foundations of Mathematics}.
\newblock Elsevier, 2003.

\bibitem{GSh93}
Dov Gabbay and Valentin Shehtman.
\newblock Undecidability of modal and intermediate first-order logics with two
  individual variables.
\newblock {\em The Journal of Symbolic Logic}, 58(3):800--823, 1993.

\bibitem{GSh98}
Dov Gabbay and Valentin Shehtman.
\newblock Products of modal logics, {P}art 1.
\newblock {\em Logic Journal of the {I}{G}{P}{L}}, 6(1):73--146, 1998.

\bibitem{GShS}
Dov Gabbay, Valentin Shehtman, and Dmitrij Skvortsov.
\newblock {\em Quantification in Nonclassical Logic, Volume 1}, volume 153 of
  {\em Studies in Logic and the Foundations of Mathematics}.
\newblock Elsevier, 2009.

\bibitem{GKV97}
Erich Gr\"{a}del, Phokion~G. Kolaitis, and Moshe~Y. Vardi.
\newblock On the decision problem for two-variable first-order logic.
\newblock {\em Bulletin of Symbolic Logic}, 3(1):53--69, 1997.

\bibitem{Halpern95}
Joseph~Y. Halpern.
\newblock The effect of bounding the number of primitive propositions and the
  depth of nesting on the complexity of modal logic.
\newblock {\em Artificial Intelligence}, 75(2):361--372, 1995.

\bibitem{HC96}
G.~E. Hughes and M.~J. Cresswell.
\newblock {\em A New Introduction to Modal Logic}.
\newblock Routledge, 1996.

\bibitem{KKZ05}
Roman Kontchakov, Agi Kurucz, and Michael Zakharyaschev.
\newblock Undecidability of first-order intuitionistic and modal logics with
  two variables.
\newblock {\em Bulletin of Symbolic Logic}, 11(3):428--438, 2005.

\bibitem{Kripke62}
Saul Kripke.
\newblock The undecidability of monadic modal quantification theory.
\newblock {\em Zeitschrift f\"{u}r Matematische Logik und Grundlagen der
  Mathematik}, 8:113--116, 1962.

\bibitem{Lowenheim15}
Leopold L\"{o}wenheim.
\newblock \"{U}ber {M}\"{o}glichkeiten im {R}elativkalk\"{u}l.
\newblock {\em Mathematische Annalen}, 76(4):447--470, 1915.

\bibitem{MMO65}
Sergei Maslov, Gregory Mints, and Vladimir Orevkov.
\newblock Unsolvability in the constructive predicate calculus of certain
  classes of formulas containing only monadic predicate variables.
\newblock {\em Soviet Mathematics Doklady}, 6:918--920, 1965.

\bibitem{Mints68}
Gregory Mints.
\newblock Some calculi of modal logic.
\newblock {\em Trudy Matematicheskogo Instituta imeni {V}. {A}. {S}teklova},
  98:88--111, 1968.
\newblock in Russian.

\bibitem{Mortimer75}
Michael Mortimer.
\newblock On languages with two variables.
\newblock {\em Zeitschrift f\"ur Mathematische Logik und Grundlagen der
  Mathematik}, pages 135--140, 1975.

\bibitem{Ono77}
Hiroakira Ono.
\newblock On some intuitionistic modal logics.
\newblock {\em Publications of the Research Institute for Mathematical
  Sciences}, 13(3):687--722, 1977.

\bibitem{Rybakov06}
Mikhail Rybakov.
\newblock Complexity of intuitionistic and {V}isser's basic and formal logics
  in finitely many variables.
\newblock In Guido Governatori, Ian~M. Hodkinson, and Yde Venema, editors, {\em
  Advances in Modal Logic 6}, pages 393--411. College Publications, 2006.

\bibitem{Rybakov08}
Mikhail Rybakov.
\newblock Complexity of intuitionistic propositional logic and its fragments.
\newblock {\em Journal of Applied Non-Classical Logics}, 18(2--3):267--292,
  2008.

\bibitem{RShIGPL18}
Mikhail Rybakov and Dmitry Shkatov.
\newblock Complexity and expressivity of propositional dynamic logics with
  finitely many variables.
\newblock {\em Logic Journal of the IGPL}, 26(5):539--547, 2018.

\bibitem{RShICTAC18}
Mikhail Rybakov and Dmitry Shkatov.
\newblock Complexity and expressivity of branching- and alternating-time
  temporal logics with finitely many variables.
\newblock In B.~Fischer B. and T.~Uustalu, editors, {\em Theoretical Aspects of
  Computing--ICTAC 2018}, volume 11187 of {\em Lecture Notes in Computer
  Science}, pages 396--414, 2018.



\bibitem{RShAiML18} Mikhail Rybakov and Dmitry Shkatov.  \newblock A
  recursively enumerable {K}ripke complete first-order logic not
  complete with respect to a first-order definable class of frames.
  \newblock In G.~Metcalfe G.~Bezhanishvili, G.~D'Agostino and
  T.~Studer, editors, {\em Advances in Modal Logic}, volume~12. pages
  531--540. College Publications, 2018.

\bibitem{RShIGPL19}
Mikhail Rybakov and Dmitry Shkatov.
\newblock Complexity of finite-variable fragments of propositional modal logics
  of symmetric frames.
\newblock {\em Logic Journal of the IGPL}, 27(1):60--68, 2019.

\bibitem{RSh19SL}
Mikhail Rybakov and Dmitry Shkatov.
\newblock Undecidability of first-order modal and intuitionistic logics with
  two variables and one monadic predicate letter.
\newblock {\em Studia Logica}, 107(4):695--717, 2019.

\bibitem{RSh20JLC}
Mikhail Rybakov and Dmitry Shkatov.
\newblock Recursive enumerability and elementary frame definability in
  predicate modal logic.
\newblock {\em Journal of Logic and Computation}, 30(2):549--560, 2020.

\bibitem{RSh20AiML}
Mikhail Rybakov and Dmitry Shkatov.
\newblock Algorithmic properties of first-order modal logics of the natural
  number line in restricted languages.
\newblock In Nicola Olivetti, Rineke Verbrugge, Sara Negri, and Gabriel Sandu,
  editors, {\em Advances in Modal Logic}, volume~13. pages
  523--539. College Publications,  2020.

\bibitem{RShJLC20a}
Mikhail Rybakov and Dmitry Shkatov.
\newblock Algorithmic properties of first-order modal logics of finite {K}ripke
  frames in restricted languages.
\newblock {\em Journal of Logic and Computation}, 30(7):1305--1329, 2020.

\bibitem{RShJLC21a}
Mikhail Rybakov and Dmitry Shkatov.
\newblock Complexity of finite-variable fragments of products with {K}.
\newblock {\em Journal of Logic and Computation}, 31(2):426--443,
2021.

\bibitem{RShJLC21b}
Mikhail Rybakov and Dmitry Shkatov.
\newblock Algorithmic properties of first-order superintuitionistic logics of
  finite {K}ripke frames in restricted languages.
\newblock {\em Journal of Logic and Computation}, 31(2):494--522, 2021.

\bibitem{Shehtman19}
Valentin Shehtman and Dmitry Shkatov.
\newblock On one-variable fragments of modal predicate logics.
\newblock In {\em Proceedings of SYSMICS2019}, pages 129--132. Institute for
  Logic, Language and Computation, University of Amsterdam, 2019.

\bibitem{ShSh20}
Valentin Shehtman and Dmitry Shkatov.
\newblock Some prospects for semiproducts and products of modal logics.
\newblock In {\em Short Papers. Advances in Modal Logic 2020}, pages 107--111.
  University of Helskinki, 2020.

\bibitem{WZ01}
Frank Wolter and Michael Zakharyaschev.
\newblock Decidable fragments of first-order modal logics.
\newblock {\em The Journal of Symbolic Logic}, 66:1415--1438, 2001.

\end{thebibliography}

\documentclass[12pt,fleqn]{article}

\usepackage{geometry}

\usepackage[cp1251]{inputenc}
\usepackage[T2A]{fontenc}
\usepackage[english]{babel}

\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{array}
\usepackage[auth-sc]{authblk}
\usepackage{bm}
\usepackage{cite}
\usepackage{color}
\usepackage{dsfont}
\usepackage{enumerate}
\usepackage{enumitem}
\usepackage{epsf}
\usepackage{euscript}
\usepackage{graphicx} 
\usepackage{indentfirst}
\usepackage{latexsym}
\usepackage{mathrsfs}
\usepackage{mathtools}
\usepackage{tikz}
\usepackage{pgf}

\newtheorem{theorem}             {Theorem}

\newlength{\templength}

\newcommand{\dis}{\vee}

\newcommand{\imp}{\rightarrow}

\begin{document}

\selectlanguage{english}

\title{Algorithmic properties of $\mathbf{QK4.3}$ and
  $\mathbf{QS4.3}$}

\author[1]{M.\,Rybakov}
\author[2]{D.\,Shkatov}
\affil[1]{IITP RAS, HSE University, Tver State University}
\affil[2]{University of the Witwatersrand, Johannesburg}

\date{} 

%
%

\maketitle

\begin{abstract}
  We prove that predicate modal logics $\mathbf{QK4.3}$ and
  $\mathbf{QS4.3}$ are undecidable---more precisely,
  $\Sigma^0_1$-complete--in languages with two individual variables,
  one modandic predicate letter, and one proposition letter.  
\end{abstract}


It is natural to expect that predicate modal logics should be
algorithmically harder than the classical predicate logic
$\mathbf{QCl}$, just as propositional modal logics are, as a rule,
algorithmically harder than the classical propositional logic.
Nevetheless, numerous predicate modal logics are just as hard as
$\mathbf{QCl}$, i.e., $\Sigma^0_1$\nobreakdash-complete: some---such
as $\mathbf{QK}$, $\mathbf{T}$, $\mathbf{S4}$, and $\mathbf{S5}$---are
recursively axiomatizable over $\mathbf{QCl}$~\cite{HC96,GShS}; others
are recursively embeddable~\cite{RShAiML18,RSh20JLC} into
$\mathbf{QCl}$ through the standard traslation.  It, however, turns
out that $\Sigma^0_1$\nobreakdash-complete predicate modal logics can
be distinguished from $\mathbf{QCl}$ by algorithmic properties of
their fragments: while the monadic fragment of $\mathbf{QCl}$ is
decidable~\cite{Lowenheim15,Behmann22}, the monadic fragments of most
$\Sigma^0_1$-complete modal logics are not~\cite{Kripke62}; while the
two-variable fragment of $\mathbf{QCl}$ is
decidable~\cite{Mortimer75,GKV97}, the two-variable fragments of most
$\Sigma^0_1$\nobreakdash-complete modal predicate logics are
not~\cite{KKZ05,RSh19SL}.  This leads to the study of the algorithmic
properties of the {\em fragments}\/ of modal predicate logics.

The algorithmic properties of one-variable and two-variable fragments
of first-order modal logics are also of interest due to close links
between those fragments and, respectively, two-dimensional and
three-dimensional propositional modal
logics~\cite{GSh98,GKWZ,Shehtman19,ShSh20,RShJLC21a}.

The study of the algorithmic properties of fragments of predicate
modal, and related superintuitionistic, logics
(see~\cite{Kripke62,MMO65,Mints68,Ono77,Gabbay81,AD90,GSh93,WZ01,KKZ05,RSh19SL,RShJLC20a,RShJLC21b};
for a summary of results, see~\cite[Introduction]{RShJLC20a}) is much
less advanced than similar research for $\mathbf{QCl}$~\cite{BGG97}.

In the present paper, we attempt to identify the minimal undecidable
fragments of the predicate counterparts $\mathbf{QK4.3}$ and
$\mathbf{QS4.3}$ of the well-know propositional modal logics
$\mathbf{K4.3}$ and $\mathbf{S4.3}$.  It is known~\cite{Corsi93} that
$\mathbf{QK4.3}$ and $\mathbf{QS4.3}$ are finitely axiomatizable over
$\mathbf{QCl}$ and, hence, they are $\Sigma^0_1$-complete.  The logics
$\mathbf{QK4.3}$ and $\mathbf{QS4.3}$ are faithfully characterized
using Kripke semantics with expanding domains~\cite{HC96}, \cite[\S
3.1]{GShS}: they are determined by all, respectively, strict and
partial linear orders~\cite{Corsi93}.  A closely related logic
$\mathbf{QK4.3.D.X}$ is determined by the rationals with the natural
strict order, viewed as a Kripke frame~\cite{Corsi93}.

The main interest of the results presented here is due the techniques
used: the known techniques for proving lower bounds in predicate modal
and superintuitionistic logics in languages with a few variables and a
few predicate letters~\cite{RSh19SL,RShJLC20a,RShJLC21b}, being based
on propositional-level
techniques~\cite{Halpern95,ChRyb03,Rybakov06,Rybakov08,RShICTAC18,RShIGPL18,RShIGPL19}
developed for logics of frames with unbounded branching, are
inapplicable to logics of linear frames (the only exception being our
earlier work~\cite{RSh20AiML}, where the techniques used for
establishing results reported here originate).

Predicate modal languages are obtained by enriching the classical
preciate language with a unary modal connective $\Box$ (for more
backgroud on predicate modal logic, see~\cite{HC96,FM98,GShS}

To recall the definitions of the logics we study, we use the following
notation: if $\Gamma$ is a set of formulas and $\varphi$ is a formula,
$\Gamma\oplus\varphi$ denotes the closure of $\Gamma\cup\{\varphi\}$
under modes ponens, generalization, necessitation, and predicate
substitution.  Then,

\settowidth{\templength}{\mbox{$\mathbf{QK4.3}$}}
$$
\begin{array}{lcl}
\mathbf{QK}
  & =
  & \parbox{\templength}{$\mathbf{QCl}$}
    ~\oplus~ \Box (p\to q) \imp (\Box p\to \Box q);
  \smallskip\\
\mathbf{QK4}
  & =
  & \parbox{\templength}{$\mathbf{QK}$}
    ~\oplus~ \Box p \imp \Box \Box p;
  \smallskip\\
\mathbf{QS4}
  & =
  & \parbox{\templength}{$\mathbf{QK4}$}
    ~\oplus~ \Box p \imp p;
  \smallskip\\
\mathbf{QS4.3}
  & =
  & \parbox{\templength}{$\mathbf{QS4}$}
    ~\oplus~ \Box (\Box p \imp q) \dis \Box (\Box q \imp p);
  \smallskip\\
\mathbf{QK4.3}
  & =
  & \parbox{\templength}{$\mathbf{QK4}$}
    ~\oplus~ \Box (p\wedge\Box p \imp q) \dis \Box (q\wedge\Box q \imp p);
  \smallskip\\
\mathbf{QK4.3.D.X}
  & =
  & \parbox{\templength}{$\mathbf{QK4.3}$}
    ~\oplus~ \Diamond \top
    ~\oplus~ \Box \Box p \imp \Box p.
\end{array}
$$



\begin{theorem}
  \label{thr:1}
  Logics $\mathbf{QK4.3}$, $\mathbf{QS4.3}$, and $\mathbf{QK4.3.D.X}$
  are $\Sigma^0_1$\nobreakdash-complete in languages containing one
  monadic predicate letter, one proposition letter, and two individual
  variables.
\end{theorem}

Theorem~\ref{thr:1} is proved as follows: we encode a
$\Sigma^0_1$\nobreakdash-hard tiling problem~\cite{Berger66} using
predicate modal formulas with only two variables, and only binary and
unary predicate letters; then, we simulate binary letters with monadic
one; finally, we simulate the monadic letters with just one monadic
and one nullary letter; the latter reductions do not use more than two
individual variables.

This result can be extended to logics containing the Barcan formula
$\mathit{bf} = \forall x\,\Box P(x)\to \Box\forall x\, P(x)$:

\begin{theorem}
  \label{thr:2}
  Logics $\mathbf{QK4.3}\oplus\mathit{bf}$,
  $\mathbf{QS4.3}\oplus\mathit{bf}$, and
  $\mathbf{QK4.3.D.X}\oplus\mathit{bf}$ are
  $\Sigma^0_1$\nobreakdash-complete in the language containing one
  monadic predicate letter, one proposition letter, and two individual
  variables.
\end{theorem}

Furthermore, we obtain the following generalisation of
Theorems~\ref{thr:1} and~\ref{thr:2}:

\begin{theorem}
  \label{thr:3}
  Every logic in the intervals
  $[\mathbf{QK4.3},\mathbf{QK4.3.D.X}\oplus\mathit{bf}]$ and
  $[\mathbf{QS4.3},\mathbf{QS4.3}\oplus\mathit{bf}]$ are
  $\Sigma^0_1$\nobreakdash-hard in the language containing one monadic
  predicate letter, one proposition letter, and two individual
  variables.
\end{theorem}

Our proofs of Theorems~\ref{thr:1}--\ref{thr:3} rely on Kripke
completeness of $\mathbf{QK4.3}$, $\mathbf{QS4.3}$, and
$\mathbf{QK4.3.D.X}$ and on soundness of
$\mathbf{QK4.3}\oplus\mathit{bf}$, $\mathbf{QS4.3}\oplus\mathit{bf}$,
and $\mathbf{QK4.3.D.X}\oplus\mathit{bf}$ with respect to,
respectively, $\mathbf{QK4.3}$-frames, $\mathbf{QS4.3}$-frames, and
$\mathbf{QK4.3.D.X}$-frames with constant domains.  

\bigskip

\textit{
  This research has been supported by the Russian Science Foundation
  with grant \mbox{21--18--00195}; it has been carried out at Tver
  State University.
}


\end{document}